OpenStudy (anonymous):

the grammy's sale cd's for 8.00 $ each + shippinq and handlinq 15.00 $ entire order ! what's the greatest number of cd's you can order with a 100.00 $ gift card . show your work ?

OpenStudy (radar):

Shipping and handling=$15.00 Amount on gift card =$100 - sh=$85 for purchase 85/8 =10 CD's

OpenStudy (anonymous):

let x= amount CDs so, x*8 = cost of all CDs Plus shipping.. +15 so, 8x+15 = 100 8x= 100-15 8x=85 x=85/8 x=has to be integer (you can't order10.6 CDs.. right?) so, you can order 10 CDs